# Adaptive Pruning for Increased Robustness and Reduced Computational Overhead in Gaussian Process Accelerated Saddle Point Searches

### v2

@@ 2.2 Optimal Transport GP
The robustness and computational scaling of GP-accelerated searches link directly to the metric used to define dissimilarity between molecular configurations, $D(\mathbf{x}_{i},\mathbf{x}_{j})$. Metrics that depend on fixed atom indices, such as the “1D max log” metric (Eq. 7), lack invariance to the permutation of identical atoms. Even low-energy events, like rotations of a methyl group, may swap indices of two hydrogen atoms, resulting in a large, unphysical distance in this metric. To solve this, we ground our dissimilarity measure in Optimal Transport (OT) theory [57, 34]. OT provides a deterministic prescription to “lift” a geometric distance between points into a distance between entire distributions of points. We treat each molecular configuration $\mathbf{x}$ as a discrete distribution of element-typed (i.e., “colored”) points in three-dimensional space. Conceptually, OT frames the distance calculation as a “transport problem.” If we consider a configuration $\mathbf{x}_{i}$ to be piles of “mass” (the atoms) and configuration $\mathbf{x}_{j}$ as a set of “holes” into which the mass must be moved, then the cost of moving one unit of mass from a point $\mathbf{r}_{a}$ to a point $\mathbf{r}_{b}$ is the Euclidean distance $\|\mathbf{r}_{a}-\mathbf{r}_{b}\|$. The Earth Mover’s Distance (EMD), the computational analogue of the Wasserstein-1 distance, finds the “transport plan” that minimizes the total work, i.e., the (mass $\times$ distance) summed over all movements. By enforcing coloring, we have that “carbon” masses only move to “carbon” holes, respecting chemical identity. More importantly, the EMD inherently solves the permutation problem as the optimal transport plan corresponds to the optimal matching between the two sets of atoms, automatically assigning atom 1 in $\mathbf{x}_{i}$ to its “correct” counterpart in $\mathbf{x}_{j}$, regardless of its index. This yields a physically meaningful distance that correctly identifies symmetrically equivalent configurations as identical. We leverage this robust, permutation-invariant metric in our Farthest Point Sampling (FPS) to select a truly diverse set of configurations for hyperparameter optimization. Figure 2 summarizes the key modifications to the algorithm workflow in this improved GP acceleration of saddle point searches.
@@ 3.1.1 Reliability
Figure 3: Comparison of the performance of the OTGPD saddle point search method against GPDimer and regular dimer without GP acceleration for 238 molecular systems. A calculation exceeding 240 minutes or raising an error in the electronic structure calculation counts as a failure. (A, B) Success outcomes for each system, ordered along the horizontal axis by mean root-mean-square deviation (RMSD) between the saddle point and the initial atomic structure. Red dots denote systems where only OTGPD succeeded, blue dots where only the alternate method succeeded, and orange dots where both failed. (C) Bar chart summarizing success rate of each method for initial structures that represent a single fragment and those representing two fragments. The single-fragment cases represent typical saddle point searches, while the two-fragment cases test performance on more complex, often arbitrary dissociation pathways. The OTGPD method demonstrates a clear advantage. In addition to a mutual success rate exceeding 93% in both comparisons, OTGPD uniquely finds the saddle point for an additional 11 systems (4.6%) compared to GPDimer (1.3%) and additional 9 systems (3.8%) compared to Dimer (1.7%).
@@
We further investigate the four systems (Fig. 3 (A), blue dots) where the GPDimer runs succeed but the OTGPD runs fail. This discrepancy does not stem from a flaw in the OTGPD algorithm but from an artifact of the dimer initialization procedure (visualized in Fig. 2). For OTGPD runs, we form the initial dimer configuration by displacing atoms along the softest mode found by an inexpensive semi-empirical (xtb, GFN2-xTB) calculation. In these specific cases, this displacement produced an unphysical starting geometry with high atomic forces or atoms in close contact. When any GP-accelerated method begins from such a high-energy, high-force baseline, the initial surrogate model learns high forces and energies corresponding to unphysical geometries are not anomalous, thus “poisoning” the run, leading to optimization failures. The GPDimer benchmark used a different initialization method which, for these systems, happened to produce a viable starting configuration. Robust dimer initialization routines are not a solved problem, nor the focus of this work. As detailed in the SI, when the OTGPD method starts from the same initial geometries used by the GPDimer, it converges successfully and more efficiently. Across these four systems, the re-run OTGPD required a near-identical total number of HF evaluations (146 for OTGPD vs. 148 for GPDimer) while completing in less than half the average wall time (OTGPD, 51.3 minutes vs. GPDimer, 109.8 minutes), despite running with half the number of parallel processes (OTGPD 8 vs GPDimer 16).
@@ 3.1.2 Saddle Consistency
Our analysis, therefore, focuses on algorithmic consistency across the full set of mutually successful runs. When compared against the baseline dimer method, the OTGPD method demonstrates high fidelity, locating transition states of effectively equal energy in 89.7% of cases. In the few instances with differing outcomes, the energy differences remain small (¡ 0.05 eV) and show no systematic bias toward either method. The comparison with GPDimer reveals a similarly high degree of agreement, with both methods finding saddle points with the same energy in over 75% of the systems.
@@ 4.3 Hyperparameter Guardrails
Embedding the logarithmic barrier of Eq. (16) directly into the objective function serves as a “hard wall” to enforce this constraint. Unlike a soft penalty, the barrier creates a diverging gradient (Eq. 18) as the optimizer approaches the bound $\lambda_{\max}$, which guarantees that the optimizer cannot enter the unstable region. Such constraints on $\sigma_{f}^{2}$ are interpretable, with a physically meaningful upper bound rather than a more abstract L1/L2 penalty strength. We found this gradient-based barrier necessary as a simple hard bound proved insufficient, with the optimizer “sticking” to the boundary.
@@ Acknowledgments
R.G. received funding from the Icelandic Research Fund (grant number 217436-053) and financial support from Dr. Guillaume Fraux and Prof. Michele Ceriotti of Lab-COSMO. The authors thank Prof. Birgir Hrafnkelsson, Prof. Thomas Bligaard, Dr. Andreas Vishart and Dr. Miha Gunde for helpful discussions. R.G. also acknowledges valuable discussions with Dr. Amrita Goswami, Dr. Moritz Sallermann, Prof. Debabrata Goswami, Mrs. Sonaly Goswami, and Mrs. Ruhila Goswami. The calculations were carried out using resources supplied by the Icelandic Research e-Infrastructure project (IREI), supported by the Icelandic Infrastructure Fund. R.G. presented a preliminary account of the work presented here at UNOOS 2025, a conference held in honor of Prof. Debabrata Goswami’s on the occation of his sixtieth birthday. R.G. warmly dedicates this article to him on this occasion.
@@
Appendix S1 Reproduction note
@@
Appendix S2 Rotation removal implementation
@@
Appendix S3 More computational details
@@
Appendix S4 Distance measure quantification
@@
This sensitivity to labeling motivates using the intensive EMD. Figure S1 demonstrates this, by contrasting the behavior of both metrics for the asymmetric stretching of a water molecule.
Figure S1: Comparison of the 1D max log distance and the Earth Mover’s Distance (EMD) for an asymmetric stretch of a water molecule. While configuration $x_{2}$ and $x_{2,s}$ are physically identical (differing only by the permutation of hydrogen atom labels), the 1D max log metric incorrectly assigns a large distance between them and the reference $x_{1}$. In contrast, the EMD correctly identifies them as being equidistant from the reference, demonstrating its permutational invariance.
S4.1 Notational clarity
@@
Appendix S5 Fragment detection
@@
From the resulting bond order matrix, a molecular connectivity graph was constructed. A chemical bond between two atoms, i and j, was defined to exist if and only if their calculated bond order, $BO_{i}j$, exceeded a threshold, 0.7 here. Based on a calibration procedure for representative configurations, a bond order threshold of 0.7 was selected. This value provides a robust margin to exclude minor, non-zero bond orders (up to $\approx 0.56$) calculated for close inter-fragment contacts, while safely including all genuine covalent bonds (typically with BOij ¿ 0.9). Finally, the distinct molecular fragments were identified by determining the connected components of the resulting graph. Fragments with geometric centers less than 1 angstrom away were merged together to take into account slightly fragmented systems.
Appendix S6 D136 and the effect of initial conditions
@@
S6.2 Post-Initialization Comparability for D136
@@
Appendix S7 Statistical analysis of cost
@@
We fit hierarchical Bayesian regression models (negative-binomial for the number of PES evaluations and Gamma for the wall-clock time) with the method (OT-GPD, GP-Dimer, standard Dimer) as a fixed effect and a random intercept for each molecule–spin combination [16]. We report posterior medians together with 95% credible intervals (CrI) in Tbl. S1.
Table S1: Table of modeled performance metrics
@@
The statistical model confirms the practical gains of the OT-GP framework. Compared with the baseline Dimer, OTGPD achieves an $\approx$ 89% reduction in the median number of expensive force evaluations (254 → 28 calls) and more than halves the median time-to-solution (20.5 min → 9.0 min). The improvement over the earlier GP-Dimer also proves decisive: the OTGPD runs $\approx$ 30% faster (9.0 min versus 12.8 min) while using slightly fewer force calls. The posterior distributions for these effects show clear separation from zero, indicating a ¿99% probability that OTGPD outperforms both comparators.
@@
S7.1 Model specifications
S7.1.1 Prior distributions
We use weakly informative priors, chosen to regularise the models without imposing strong beliefs about the magnitudes of effects, detailed in Table S2.
Table S2: Priors for predictive model
@@
[…]
@@
With the exact results in Table S3.
Table S3: Results of the PES model
@@
S7.1.4 Total wall-clock time
@@
We show the model predictions overlaid with the data in S2.
Figure S2: Posterior predictions from the hierarchical Gamma spline model for total wall-clock time as a function of PES evaluations. Points represent the raw data for each of the three methods. Solid lines are the model’s posterior mean predictions, with shaded regions indicating the 95% credible intervals. The relationship between time and PES calls is allowed to be non-linear and is modeled independently for each method. For the GP-based methods, dashed lines indicate where the model is extrapolating beyond the maximum number of observed PES calls for that method, which are marked by the vertical dashed lines.
@@
Appendix S8 GPDimer / OTGPD failure modes
@@
When any GP-accelerated method is “poisoned” with such a high-energy, high-force baseline, the initial surrogate model learns that these unphysical configurations are normal, leading to an unstable search that fails to converge. The GPDimer benchmark, used a different (and in these cases, fortuitously non-pathological) initialization based on displacing the least coordinated atom, bypassed this specific problem. Figure S3 shows the initial points at the start of the algorithm runs, extracted from the trajectory files.
Figure S3: Initializations for the GPDimer (top) and OTGPD (bottom). The XTB initialization procedure in these four cases results in starting the runs from unphysical initial geometries leading to failures in the optimization routine. With XTB, D016 has unphysically close carbon atoms, D084 has a shortened carbon-oxygen bond, D100 has near-overlapping carbon atoms, and S242 has a CH3 end-group instead of the hydrogen on the middle carbon.
@@
Times reported in minutes –¿ total seconds / 60.
S8.1 D016

### v3

@@ 2.2 Optimal Transport GP
The robustness and computational scaling of GP-accelerated searches link directly to the metric used to define dissimilarity between molecular configurations, $D(\mathbf{x}_{i},\mathbf{x}_{j})$. Metrics that depend on fixed atom indices, such as the “1D max log” metric (Eq. 7), lack invariance to the permutation of identical atoms. Even low-energy events, such as the rotations of a methyl group, may swap indices of two hydrogen atoms, resulting in a large, unphysical distance in this metric. To solve this, we ground our dissimilarity measure in Optimal Transport (OT) theory [59, 35]. OT provides a deterministic prescription to “lift” a geometric distance between points into a distance between entire distributions of points. We treat each molecular configuration $\mathbf{x}$ as a discrete distribution of element-typed (i.e., “colored”) points in three-dimensional space. Conceptually, OT frames the distance calculation as a “transport problem.” If we consider a configuration $\mathbf{x}_{i}$ to be piles of “mass” (the atoms) and configuration $\mathbf{x}_{j}$ as a set of “holes” into which the mass must be moved, then the cost of moving one unit of mass from a point $\mathbf{r}_{a}$ to a point $\mathbf{r}_{b}$ is the Euclidean distance $\|\mathbf{r}_{a}-\mathbf{r}_{b}\|$. The Earth Mover’s Distance (EMD), the computational analogue of the Wasserstein-1 distance, finds the “transport plan” that minimizes the total work, i.e., the (mass $\times$ distance) summed over all movements. By enforcing coloring, we have that “carbon” masses only move to “carbon” holes, respecting chemical identity. More importantly, the EMD inherently solves the permutation problem as the optimal transport plan corresponds to the optimal matching between the two sets of atoms, automatically assigning atom 1 in $\mathbf{x}_{i}$ to its “correct” counterpart in $\mathbf{x}_{j}$, regardless of its index. This yields a physically meaningful distance that correctly identifies symmetrically equivalent configurations as identical. We leverage this robust, permutation-invariant metric in our Farthest Point Sampling (FPS) to select a truly diverse set of configurations for hyperparameter optimization. Figure 2 summarizes the key modifications to the algorithm workflow in this improved GP acceleration of saddle point searches.
@@ 3.1.1 Reliability
Figure 3: Comparison of the performance of the OTGPD saddle point search method against GPDimer and regular dimer without GP acceleration for 238 molecular systems. A calculation exceeding 4 hours or raising an error in the electronic structure calculation counts as a failure. (A, B) Success outcomes for each system, ordered along the horizontal axis by mean root-mean-square deviation (RMSD) between the saddle point and the initial atomic structure. Red dots denote systems where only OTGPD succeeded, blue dots where only the alternate method succeeded, and orange dots where both failed. (C) Bar chart summarizing success rate of each method for initial structures that represent a single fragment and those representing two fragments. The single-fragment cases represent typical saddle point searches, while the two-fragment cases test performance on more complex, often arbitrary dissociation pathways. The OTGPD method demonstrates a clear advantage. In addition to a mutual success rate exceeding 93% in both comparisons, OTGPD uniquely finds the saddle point for an additional 11 systems (4.6%) compared to GPDimer (1.3%) and additional 9 systems (3.8%) compared to Dimer (1.7%).
@@
We further investigate the four systems (Fig. 3 (A), blue dots) where the GPDimer runs succeed but the OTGPD runs fail. This discrepancy does not stem from a flaw in the OTGPD algorithm but from an artifact of the dimer initialization procedure (visualized in Fig. 2). For OTGPD runs, we form the initial dimer configuration by displacing atoms along the softest mode found by an inexpensive semi-empirical (xtb, GFN2-xTB) calculation. In these specific cases, this displacement produced an unphysical starting geometry with high atomic forces or atoms in close contact. When any GP-accelerated method begins from such a high-energy, high-force baseline, the initial surrogate model learns high forces and energies corresponding to unphysical geometries are not anomalous, thus “poisoning” the run, leading to optimization failures. The GPDimer benchmark used a different initialization method which, for these systems, happened to produce a viable starting configuration. Robust dimer initialization routines are not a solved problem, nor the focus of this work. As detailed in the SI, when the OTGPD method starts from the same initial geometries used by the GPDimer, it converges successfully and more efficiently. For these four systems, the re-run OTGPD required a near-identical total number of HF evaluations (146 for OTGPD vs. 148 for GPDimer) while completing in less than half the average wall time (OTGPD, 51.3 minutes vs. GPDimer, 109.8 minutes), despite running with half the number of parallel processes (OTGPD 8 vs GPDimer 16).
@@ 3.1.2 Saddle Consistency
Our analysis, therefore, focuses on algorithmic consistency across the full set of mutually successful runs. When compared against the baseline dimer method, the OTGPD method demonstrates high fidelity, locating transition states of effectively equal energy in 89.7% of cases. In the few instances with differing outcomes, the energy differences remain small (< 0.05 eV) and show no systematic bias toward either method. The comparison with GPDimer reveals a similarly high degree of agreement, with both methods finding saddle points with the same energy in over 75% of the systems.
@@ 4.3 Hyperparameter Guardrails
This specific failure mode requires a solution distinct from conventional L1 or L2 regularization [28]. Such methods apply a soft penalty to the objective to combat general overfitting, but a strong MLL gradient can overcome this penalty, failing to prevent the variance explosion seen in Fig. 4. We therefore draw inspiration from interior point methods [51] used in constrained optimization.
Embedding the logarithmic barrier of Eq. (16) directly into the objective function serves as a “hard wall” to enforce this constraint. Unlike a soft penalty, the barrier creates a diverging gradient (Eq. 18) as the optimizer approaches the bound $\lambda_{\max}$, which guarantees that the optimizer cannot enter the unstable region. Such constraints on $\sigma_{f}^{2}$ are interpretable, with a physically meaningful upper bound rather than a more abstract L1/L2 penalty strength. We find this gradient-based barrier necessary as a simple hard bound proved insufficient, with the optimizer “sticking” to the boundary.
@@ Acknowledgments
R.G. received funding from the Icelandic Research Fund (grant number 217436-053) and financial support from Dr. Guillaume Fraux and Prof. Michele Ceriotti of Lab-COSMO. The authors thank Prof. Birgir Hrafnkelsson, Prof. Thomas Bligaard, Dr. Andreas Vishart and Dr. Miha Gunde for helpful discussions. R.G. also acknowledges valuable discussions with Dr. Amrita Goswami, Dr. Moritz Sallermann, Prof. Debabrata Goswami, Mrs. Sonaly Goswami, and Mrs. Ruhila Goswami. The calculations were carried out using resources supplied by the Icelandic Research e-Infrastructure project (IREI), supported by the Icelandic Infrastructure Fund. R.G. presented a preliminary account of the work presented here at UNOOS 2025, a conference held in honor of Prof. Debabrata Goswami’s on the occasion of his sixtieth birthday. R.G. warmly dedicates this article to him on this occasion.
@@
Appendix A Reproduction note
@@
Appendix B Rotation removal implementation
@@
Appendix C More computational details
@@
Appendix D Distance measure quantification
@@
This sensitivity to labeling motivates using the intensive EMD. Figure 8 demonstrates this, by contrasting the behavior of both metrics for the asymmetric stretching of a water molecule.
Figure 8: Comparison of the 1D max log distance and the Earth Mover’s Distance (EMD) for an asymmetric stretch of a water molecule. While configuration $x_{2}$ and $x_{2,s}$ are physically identical (differing only by the permutation of hydrogen atom labels), the 1D max log metric incorrectly assigns a large distance between them and the reference $x_{1}$. In contrast, the EMD correctly identifies them as being equidistant from the reference, demonstrating its permutational invariance.
D.1 Notational clarity
@@
Appendix E Fragment detection
@@
From the resulting bond order matrix, a molecular connectivity graph was constructed. A chemical bond between two atoms, i and j, was defined to exist if and only if their calculated bond order, $BO_{i}j$, exceeded a threshold, 0.7 here. Based on a calibration procedure for representative configurations, a bond order threshold of 0.7 was selected. This value provides a robust margin to exclude minor, non-zero bond orders (up to $\approx 0.56$) calculated for close inter-fragment contacts, while safely including all genuine covalent bonds (typically with BOij > 0.9). Finally, the distinct molecular fragments were identified by determining the connected components of the resulting graph. Fragments with geometric centers less than 1 angstrom away were merged together to take into account slightly fragmented systems.
Appendix F D136 and the effect of initial conditions
@@ F.1 Initialization
F.2 Post-Initialization Comparability for D136
@@
Appendix G Statistical analysis of cost
@@
We fit hierarchical Bayesian regression models (negative-binomial for the number of PES evaluations and Gamma for the wall-clock time) with the method (OT-GPD, GP-Dimer, standard Dimer) as a fixed effect and a random intercept for each molecule–spin combination [16]. We report posterior medians together with 95% credible intervals (CrI) in Tbl. 1.
Table 1: Table of modeled performance metrics
@@
The statistical model confirms the practical gains of the OT-GP framework. Compared with the baseline Dimer, OTGPD achieves an $\approx$ 89% reduction in the median number of expensive force evaluations (254 → 28 calls) and more than halves the median time-to-solution (20.5 min → 9.0 min). The improvement over the earlier GP-Dimer also proves decisive: the OTGPD runs $\approx$ 30% faster (9.0 min versus 12.8 min) while using slightly fewer force calls. The posterior distributions for these effects show clear separation from zero, indicating a >99% probability that OTGPD outperforms both comparators.
@@
G.1 Model specifications
G.1.1 Prior distributions
We use weakly informative priors, chosen to regularise the models without imposing strong beliefs about the magnitudes of effects, detailed in Table 2.
Table 2: Priors for predictive model
@@ G.1.3 Number of PES calls
  • Fixed effects ($\beta_{method}$) encode the multiplicative shift relative to the baseline Dimer (reference level).
@@
With the exact results in Table 3.
Table 3: Results of the PES model
@@
G.1.4 Total wall-clock time
@@
We show the model predictions overlaid with the data in Figure 9.
Figure 9: Posterior predictions from the hierarchical Gamma spline model for total wall-clock time as a function of PES evaluations. Points represent the raw data for each of the three methods. Solid lines are the model’s posterior mean predictions, with shaded regions indicating the 95% credible intervals. The relationship between time and PES calls is allowed to be non-linear and is modeled independently for each method. For the GP-based methods, dashed lines indicate where the model is extrapolating beyond the maximum number of observed PES calls for that method, which are marked by the vertical dashed lines.
@@
Appendix H GPDimer / OTGPD failure modes
@@
When any GP-accelerated method is “poisoned” with such a high-energy, high-force baseline, the initial surrogate model learns that these unphysical configurations are normal, leading to an unstable search that fails to converge. The GPDimer benchmark, used a different (and in these cases, fortuitously non-pathological) initialization based on displacing the least coordinated atom, bypassed this specific problem. Figure 10 shows the initial points at the start of the algorithm runs, extracted from the trajectory files.
Figure 10: Initializations for the GPDimer (top) and OTGPD (bottom). The XTB initialization procedure in these four cases results in starting the runs from unphysical initial geometries leading to failures in the optimization routine. With XTB, D016 has unphysically close carbon atoms, D084 has a shortened carbon-oxygen bond, D100 has near-overlapping carbon atoms, and S242 has a CH3 end-group instead of the hydrogen on the middle carbon.
@@
Times reported in minutes –> total seconds / 60. Calls represent the number of Hartree-Fock calculations, which are the “true energy samples” required.
H.1 D016